### v1

# Collective optical properties of moiré excitons

Abstract
We propose that excitons in moiré transition metal dichalcogenide bilayers offer a promising platform for investigating collective radiative properties. While some of these optical properties resemble those of cold atom arrays, moiré excitons extend to the deep subwavelength limit, beyond the reach of current optical lattice experiments. Remarkably, we show that the collective optical properties can be exploited to probe certain correlated electron states. Specifically, we illustrate that the Wigner crystal states of electrons doped into these bilayers act as an emergent periodic potential for excitons. Moreover, the collective dissipative excitonic bands and their associated Berry curvature can reveal various charge orders that emerge at the corresponding electronic doping. Our study provides a promising pathway for future research on the interplay between collective effects and strong correlations involving moiré excitons.
Introduction. — The collective behavior of optical emitters in two-dimensional lattices driven by dipole-dipole interaction has generated considerable interest. Emitters in such arrays display a range of intriguing properties, including perfect reflection and transmission, dressed lineshift, enhanced and suppressed radiation, and even topological attributes such as collective Chern bands and protected edge states [1, 2, 3, 4]. Recently, this system has been experimentally realized using cold atoms in optical lattices [5], providing a promising avenue for the study of these coherent phenomena.
On the other hand, excitons (electron-hole bound states) in transition metal dichalcogenide (TMD) bilayers [6, 7, 8, 9, 10, 11] can form in a moiré lattice, providing an interesting platform to study collective optical properties. Moreover, the bilayers can host Wigner crystal (WC) states of doped electrons  [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], which can modify the excitonic properties.
In this work, we investigate the collective optical properties of moiré excitons in TMDs at incompressible states of doped electrons for various fractional fillings $\nu_{e}$. In these cases, the localized charges provide an additional superpotential, which effectively modifies the lattice for excitons (see Fig. 1) and their collective optical properties. Consequently, distinct spectral and topological properties of the excitons emerge at different WCs, offering optical signatures of these charge orders (see Table 1). Notably, such filling fractions can be easily accessed in gate-tunable experiments. Such tunability, combined with adjustable twist angles, further allows for the exploration of subwavelength emitter-array physics within different lattice configurations.
Figure 1: (a) Illustration of tunneling driven by dipole-dipole interaction (green arrow) of an exciton, as indicated by a pair of red and light blue dots, in a moiré potential (yellow) with period $a_{M}$ in the absence of doped electrons ($\nu_{e}=0$). (b) At certain finite dopings (here $\nu_{e}=\frac{1}{3}$), doped electrons (dark blue dots) form a Wigner crystal state. Due to exciton-electron repulsion, the exciton cannot populate the sites occupied by electrons (indicated by the crossed-out path). The remaining available sites form the emergent lattice.
[…]
Emergent arrays $\nu_{e}$ Superradiant $\Lambda$ $C_{\mathrm{LC}}$
Triangular $0,\frac{2}{3},\frac{3}{4},\frac{6}{7}$ $0,1$ $-1,1$
Rectangular $\frac{1}{2}$ $0,1$ $0,0$
Honeycomb $\frac{1}{3}$ $0,1$ $-1,1$
Kagome $\frac{1}{4}$ $0,1,3,4$ $-1,1,1,-1$
Table 1: Emergent lattices and collective emission properties of interlayer excitons at various filling fractions ($\nu_{e}$), which form complementary lattices of charge orders in zero-twist WSe2/WS2 [14, 25]. The third column lists the collective band indices (in energy order) $\Lambda$ with enhanced radiative decay (within the light cone) compared to the bare dipole transition rate $\gamma$. $C_{\mathrm{LC}}$ denotes the corresponding Berry curvature, summed over the light cone, upon an out-of-plane magnetic field with Zeeman splitting $\mu_{B}B=20\gamma$.
Two-band model. — We consider the lowest conduction (CB) and the highest valence (VB) bands of two TMD monolayers. Each band is labeled by the valley pseudospin, which is locked to the real spin degrees of freedom for energy scales lower than the corresponding spin-orbit splittings [26]. Stacking these layers with a tunable twist angle or (for heterobilayers) lattice mismatch generally leads to interlayer coupling with an enlarged spatial periodicity $a_{M}$. We specifically consider heterobilayers [27], as the band offsets therein allow us to capture the effect of interlayer tunneling by emergent superlattice potentials, which split the original electronic dispersion into moiré bands  [^2] Such band misalignment also suppresses the topological effects favored by layer-degenerate states [67], which is not the main focus of our work.. In addition, we consider doping electrons into the first conduction moiré band and focus on the regime with dilute electron-hole pairs [6, 29], generated via weak optical excitation. Note that within such a dilute regime, the emission properties are captured by the single excitation subspace, which can be described by a non-Hermitian Hamiltonian  [^3] In this case, it is equivalent to the full Lindbladian dynamics. While our formalism can be straightforwardly generalized, for simplicity, all (doped and optically excited) electrons are assumed to be in the same layer.
We start with the following two-band model in the absence of an external drive (see its microscopic origin in Ref. [31]):
^$\hat{\mathcal{H}}=\sum_{n=c,v}\sum_{\tau}\int d^{2}\bm{s}\hat{\psi}_{\tau}^{(n)\dagger}(\bm{s})h_{n}(\bm{s})\hat{\psi}_{\tau}^{(n)}(\bm{s})+\hat{\mathcal{V}}_{e}+\hat{\mathcal{V}}_{d}.$ (1)
[…]
$\hat{\mathcal{V}}_{d}=\frac{\omega_{p}^{2}}{c^{2}\epsilon_{0}}\int_{\bm{s},\bm{s}^{\prime}}\hat{\bm{P}}^{\dagger}(\bm{s})\cdot\bm{\mathcal{G}}\left(\omega_{p},\bm{s}-\bm{s}^{\prime}\right)\cdot\hat{\bm{P}}(\bm{s}^{\prime}),$ (2)
where $c$ is the speed of light and $\epsilon_{0}$ is the vacuum electric permittivity  [^6] Dielectric screening is neglected at the target frequency for simplicity.. $\hat{\bm{P}}^{\dagger}(\bm{s})=\sum_{\tau}\bm{d}_{\tau}^{cv}\hat{\psi}_{\tau}^{(c)\dagger}(\bm{s})\hat{\psi}_{\tau}^{(v)\dagger}(\bm{s})$ is the pair creation operator  [^7] R-stacked bilayers are considered such that the two charges have the same $\tau$ [68]., where $\bm{d}_{\tau}^{cv}$ is the transition dipole matrix element between the CB and VB at valley $\tau$. $\bm{\mathcal{G}}\left(\omega_{p},\bm{s}-\bm{s}^{\prime}\right)$ is the dyadic Green’s tensor evaluated at the frequency of the target pair state $\omega_{p}$ [31]. In the following, we use the two-band model $\hat{\mathcal{H}}$ to construct the superlattice Hamiltonian for optical excitations in the lowest energy manifold and doped electrons.
Zero doping. — At $\nu_{e}=0$, optical excitations from Eq. (1) are given by excitons. To capture the lowest composite particle, we employ the tight-binding approximation to the two-particle Hamiltonian operator from $\hat{\mathcal{H}}$, whose eigenfunctions in the tight-binding limit can be approximated by moiré-Wannier orbitals [31]. The lowest energy orbital at each supersite $\bm{R}$, denoted as $w_{\bm{R}}(\bm{s}_{c},\bm{s}_{v})$ with $\bm{s}_{c}$ and $\bm{s}_{v}$ being the coordinates of the electron and hole, respectively, defines the corresponding exciton creation operator:
$\hat{x}_{\bm{R},\tau}^{\dagger}=\int d^{2}\bm{s}_{c}d^{2}\bm{s}_{v}w_{\bm{R}}(\bm{s}_{c},\bm{s}_{v})\hat{\psi}_{\tau}^{(c)\dagger}(\bm{s}_{c})\hat{\psi}_{\tau}^{(v)\dagger}(\bm{s}_{v}).$ (3)
Projecting $\hat{\mathcal{H}}$ onto these basis states, the superlattice Hamiltonian becomes  [^8] Here, we drop the dipole-dipole interaction from the small out-of-plane transition dipole (not the permanent dipole) of interlayer excitons [43], which can be further suppressed via hybridization with intralayer excitons [69].:
$\hat{H}=\left(\omega_{\mathrm{ex}}-\frac{i\gamma}{2}\right)\hat{n}_{\mathrm{ex}}-\sum_{\bm{R}\neq\bm{R}^{\prime}}\sum_{\tau,\tau^{\prime}}t_{\bm{R},\bm{R}^{\prime}}^{\tau,\tau^{\prime}}\hat{x}_{\bm{R},\tau}^{\dagger}\hat{x}_{\bm{R}^{\prime},\tau^{\prime}},$ (4)
where $\hat{n}_{\mathrm{ex}}=\sum_{\bm{R},\tau}\hat{x}_{\bm{R},\tau}^{\dagger}\hat{x}_{\bm{R},\tau}$. $\omega_{\mathrm{ex}}$ and $\gamma$ denote the exciton frequency and decay rate (we set $\hbar=1$ hereafter), respectively, in a unit supercell problem, whereas the true excitation spectrum in the superlattice is renormalized by the tunneling:
$t_{\bm{R},\bm{R}^{\prime}}^{\tau,\tau^{\prime}}=t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}\delta_{\tau,\tau^{\prime}}-\frac{\omega_{\mathrm{ex}}^{2}|d|^{2}}{c^{2}\epsilon_{0}}\bm{e}_{\tau}^{\ast}\cdot\bm{\mathcal{G}}^{\mathrm{ex}}\left(\bm{R}-\bm{R}^{\prime}\right)\cdot\bm{e}_{\tau^{\prime}},$ (5)
which incorporates the sector from charge dynamics (denoted as $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}$) and the dipole-dipole interaction from the in-plane components of $\bm{d}_{\tau}^{cv}$  [^9] Similar interactions between dipoles within the same supercell has already been incorporated in $\omega_{\mathrm{ex}}$ and $\gamma$., denoted as $d\bm{e}_{\tau}^{\ast}$ with $\bm{e}_{\tau}=\frac{\bm{e}_{x}+i\tau\bm{e}_{y}}{\sqrt{2}}$ ($\bm{e}_{x}$ and $\bm{e}_{y}$ are in-plane unit vectors) [42, 43]. Here, $\bm{\mathcal{G}}^{\mathrm{ex}}\left(\bm{R}-\bm{R}^{\prime}\right)=\int_{\bm{s},\bm{s}^{\prime}}w_{\bm{R}}^{\ast}(\bm{s})\bm{\mathcal{G}}\left(\omega_{\mathrm{ex}};\bm{s}-\bm{s}^{\prime}\right)w_{\bm{R}^{\prime}}(\bm{s}^{\prime})$ where $w_{\bm{R}}(\bm{s})\equiv w_{\bm{R}}(\bm{s},\bm{s})$ is approximated as a Gaussian with width $a_{W}$ for simplicity. We refer to the supplementary material for details of $\hat{H}$ [31].
Finite doping with charge order. — We can generalize Eq. (4) to finite $\nu_{e}$, where doped electrons form WCs. These charge orders could emerge from a generalized Hubbard model [44, 45], which naturally appears by projecting Eq. (1) onto the first moiré band of doped charges [31]. Assuming that the electronic states being stable against the dynamics of dilute optical excitations, we can treat the WC as a spatially periodic detuning with strength characterized by $\omega_{t}-\omega_{\mathrm{ex}}$, where $\omega_{t}$ is the energy of the lowest three-body state (one doped electron together with the excited electron-hole pair) within a supercell. In particular, in the regime $|\omega_{t}-\omega_{\mathrm{ex}}|\gg|t_{\bm{R},\bm{R}^{\prime}}^{\tau,\tau^{\prime}}|$, the low-energy optical excitations are either within the $\sim\omega_{\mathrm{ex}}$ or $\sim\omega_{t}$ manifold, depending on the sign of $\omega_{t}-\omega_{\mathrm{ex}}$. We focus specifically on the case $\omega_{\mathrm{ex}}\ll\omega_{t}$, which is realizable by interlayer excitons [31], to study the collective behavior of these emitters. In this subspace, these excitons are still described by the Hamiltonian Eq. (4101010Note that the energy integrals are different for the excitations in the $\sim\omega_{\mathrm{ex}}$ and $\sim\omega_{t}$ manifolds, due to their distinction in wavefunctions, except that now $\{\bm{R}\}$ lies in the emergent lattices, which are complementary lattices of the WCs. Accordingly, different emergent lattices can be realized simply by accessing distinct WCs in moiré TMD bilayers, which can be tuned by the electron filling fraction $\nu_{e}$ (doped electrons per supercell) [44, 14, 25]. We list a few possible emergent lattices observed in WSe2/WS2 in Table 1.
Figure 2: Collective excitonic lineshifts $\Delta_{\bm{Q}}$ (solid lines) and linewidths $\Gamma_{\bm{Q}}$ (dashed) emerging from charge-ordered zero-twist WSe2/WS2 with electron fillings $\nu_{e}=0$ (a) and $\nu_{e}=\frac{1}{3}$ (b). The vertical axes are displayed in units of $\gamma N_{\nu_{e}}$, with $N_{\nu_{e}}$ defined in Eq. (6). The horizontal axes show Bloch momenta $\bm{Q}$ at high symmetry points, which follow a piecewise-linear path in the Brillouin zone, as depicted by the dashed hexagon in the inset of (a). Momenta within the light cone are indicated by the green shaded area (size enlarged for clarity). Different colors label distinct single-particle exciton bands. The parameters used are: $a_{M}=8.25$nm, $\omega_{\mathrm{ex}}=1.55$eV, and $a_{W}=2$nm [31, 18].
Collective bands. — We proceed to study the Liouvillian spectrum incorporating tunneling driven by coherent and incoherent dipole-dipole interactions. To capture the relevant physics, we set $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}=0$ in Eq. (5) throughout this work  [^11] Neglecting $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}$ is a valid approximation at large $a_{M}$ because it decays exponentially with $|\bm{R}-\bm{R}|/a_{M}$, whereas the dipole-dipole interaction only scales as a power law.. Diagonalization of the Hamiltonian Eq. (4) yields the emitter spectrum $\omega_{\mathrm{ex}}+\Delta_{\bm{Q}}-\frac{i}{2}\Gamma_{\bm{Q}}$ characterized by the center-of-mass Bloch momentum $\bm{Q}$. Note that, if the excitonic states at all $\bm{R}$ constructively interfere, $\Gamma_{\bm{Q}}$ should scale with $\gamma N_{\nu_{e}}(\theta)$ [1], where:
$N_{\nu_{e}}(\theta)=(1+\theta^{2}/\delta^{2})N_{\nu_{e}},\quad N_{\nu_{e}}=\lambda_{\mathrm{ex}}^{2}/\mathcal{A}_{\nu_{e}},$ (6)
with $\mathcal{A}_{\nu_{e}}$ denoting the emergent unit cell area at electron filling $\nu_{e}$ and zero twist, $\delta$ defining the lattice mismatch between the two monolayers [18]. As such, we present the collective spectrum in units of $\gamma N_{\nu_{e}}(\theta)$ hereafter to indicate the extent of constructive interference of the exciton eigenstates.
In Fig. 2, we illustrate the collective lineshifts $\Delta_{\bm{Q}}$ and linewidths $\Gamma_{\bm{Q}}$ for the untwisted WSe2/WS2 bilayer at $\nu_{e}=0$ and $\frac{1}{3}$, corresponding to triangular and honeycomb lattices, respectively. The cases with Kagome and rectangular lattices can be found in the supplementary materials [31]. For all $\nu_{e}$ of interest, both quantities vary on scales of $N_{\nu_{e}}\gamma$ with $N_{\nu_{e}}\sim 10^{4}$. Qualitatively, this is because emissions from supersites much closer than the exciton wavelength $\lambda_{\mathrm{ex}}$ can constructively interfere.
The collective behavior is qualitatively distinct within and outside of the LC. For all $\nu_{e}$, $\Gamma_{\bm{Q}}$ is significantly greater than $\gamma$ within the LC but is suppressed outside, consistent with the fact that only emitters in the LC couple to light and radiate due to momentum conservation. Note also that such a large decay rate in the LC indicates that an emitter tends to radiate before it hops to other supersites, which is also reflected by the relatively flat $\Delta_{\bm{Q}}$ (with respect to $\bm{Q}$) therein.
Figure 3: Dependence of collective lineshifts and linewidths at $\bm{Q}=0$ (a) on $\nu_{e}$ for all states at zero twist and (b) on twisting angle $\theta$ and $\nu_{e}$, whose effects manifest in the combined factor $N_{\nu_{e}}(\theta)$, see Eq. (6). Dashed lines in (b) indicate linear fits with slope $1$. Here we set $\delta=0.04$ [18], and $a_{W}(\theta)=(1+\theta^{2}/\delta^{2})^{-\frac{1}{4}}a_{W}(0)$ [48]; all other parameters are the same as in Fig. 2.
Aside from these common properties, collective bands exhibit several distinct features at different $\nu_{e}$. We specifically illustrate how $\Delta_{0}$ and $\Gamma_{0}$ vary with $\nu_{e}$ in Fig. 3(a), as these quantities are directly accessible via reflection measurements at normal incidence [1]. First, different emergent lattices provide a distinct number of levels  [^12] Although the number of collective bands should equal two (coupled valleys) times the number of sublattices per unit emergent cell, we note that some states are degenerate and non-resolvable at high symmetry momenta.. Although this serves as the most direct signature to distinguish emergent lattices at each $\nu_{e}$, we note that states other than the lowest one possess a suppressed decay rate (except at $\nu_{e}=\frac{1}{4}$ and $\frac{1}{2}$), such that we expect they barely contribute to optical measurements. In contrast, the lowest state at each $\nu_{e}$, which is typically the most radiative one, could respond nontrivially to optical probe light; their collective properties — the overall increasing $\Delta_{0}$ and decreasing $\Gamma_{0}$ with $\nu_{e}$ — could serve as characteristics of the corresponding emergent lattices.
Finally, these spectral properties can also be tuned via twisting angle $\theta$. Fig. 3(b) illustrates the dependence of $\Delta_{0}$ and $\Gamma_{0}$ on $N_{\nu_{e}}(\theta)$. Crucially, for the same type of lattice geometry, the collective linewidths scale inversely with the emergent unit cell area, which can be confirmed by the analytic expression of the Green’s tensor at $\bm{Q}=0$. These scalings provide an indirect probe for the emergent lattice geometry: for instance, all triangular arrays (e.g., $\nu_{e}=0,\frac{2}{3}$) belong to the same line in Fig. 3(b).
Berry curvature. — Topological features can also emerge in these collective excitonic bands. To see this, we break time-reversal symmetry by adding an out-of-plane external magnetic field $B$, which (together with the magnetic dipole $\mu_{B}$) introduces a Zeeman splitting $\hat{H}\to\hat{H}^{\prime}\equiv\hat{H}+\mu_{B}B\sum_{\tau=\pm,\bm{R}}\tau\hat{x}_{\bm{R},\tau}^{\dagger}\hat{x}_{\bm{R},\tau}$ between the valley-degenerate doublets, and compute the Berry curvature $\Omega({\bm{Q}})=i\nabla_{\bm{Q}}\times\langle\bm{Q}|\nabla_{\bm{Q}}|\bm{Q}\rangle$ for each band at several $\nu_{e}$, where $|\bm{Q}\rangle$ are right eigenstates of $\hat{H}^{\prime}$  [^13] Here, without loss of generality, we choose the right eigenvectors of the non-Hermitian Hamiltonian to define the Berry curvature [70, 71]. We leave a rigorous study of non-Hermitian topology for future work. We specifically focus on $\Omega({\bm{Q}})$ of (directly) optically accessible states (i.e., within the LC) and compute the numerical results for $\nu_{e}=0$ and $\frac{1}{2}$ in Fig. 4(a), evaluated under a standard momentum discretization scheme [51].
At $\nu_{e}=0$, both collective excitonic states (labeled with respect to their energies by $\Lambda=0,1$) exhibit nontrivial and opposite $\Omega({\bm{Q}})$ centered at $\bm{Q}=0$. Notably, the summation of $\Omega({\bm{Q}})$ in LC, $C_{\mathrm{LC}}$, for these bands takes integer values $\mp 1$. Such a property results from the phase winding of the light-mediated coupling between the valley-doublets, which appears as the off-diagonal terms in the effective low momentum approximation of $\hat{H}^{\prime}$, denoted as $\hat{h}(\bm{Q})$ [31]. With $\bm{Q}=Q(\cos\phi\bm{e}_{x}+\sin\phi\bm{e}_{y})$, where $\phi$ is the polar angle of $\bm{Q}$, we have:
$\hat{h}(\bm{Q})\simeq\omega_{\mathrm{ex}}+\Delta_{0}-\frac{i\Gamma_{0}}{2}+\begin{bmatrix}\mu_{B}B&iJQ^{2}e^{-2i\phi}\\ iJQ^{2}e^{2i\phi}&-\mu_{B}B\end{bmatrix},$ (7)
where $J$ is a (generally complex) coefficient characterizing the band curvature around $\bm{Q}=0$. We find that Eq. (7) can approximately capture $\Omega(\bm{Q})$, particularly the sharp feature at $Q=\sqrt{|\mu_{B}B/J|}$.
In contrast, $C_{\mathrm{LC}}$ from generic arrays (see table 1) may be trivial if the low-$Q$ model of the target band is different from Eq. (7). For instance, both collective excitonic bands from the rectangular lattice at $\nu_{e}=\frac{1}{2}$ give zero Berry curvature, as shown in Fig. 4(a), and hence $C_{\mathrm{LC}}=0$. This is due to the absence of $C_{3}$ rotational symmetry, leading to nonzero off-diagonal terms in the corresponding Hamiltonian at $Q=0$, which indicates a gap separating two topologically trivial bands [31].
Figure 4: Comparison between properties of topological and non-topological collective bands from $\nu_{e}=0$ and $\nu_{e}=\frac{1}{2}$, respectively, with Zeeman splitting $\mu_{B}B=20\gamma$. (a) Berry curvature ($\Omega$) distribution in Bloch momentum magnitude $Q$ for the two bands at each $\nu_{e}$ (labeled by colors and $\Lambda$), with solid and dashed lines indicating results from the full model $\hat{H}^{\prime}$ and the low-momentum model Eq. (7), respectively. The vertical axis is in units of $J/\gamma$, where $J$ is fit from $\Gamma_{\bm{Q}}$ in Fig. 2(a) using a momentum sample within $Q\leq{\pi}/{25\lambda_{\mathrm{ex}}}$. The gray vertical line indicates the momentum satisfying $JQ^{2}=\mu_{B}B$. The inset shows that the dependence of $\Omega$ on the polar angle of $\bm{Q}$ is negligible. (b) The phase of $S_{-+}$ at $k=\frac{2\pi}{\lambda_{\mathrm{ex}}}$ and $\bm{k}_{||}=k_{x}\bm{e}_{x}+k_{y}\bm{e}_{y}$ for each $\nu_{e}$. All other parameters are the same as in Fig. 2.
Similar arguments also apply to Berry curvatures of higher collective states. More specifically, the suppressed $J$ for $\Lambda=2,3$ at $\nu_{e}=\frac{1}{3}$ and $\Lambda=2,5$ at $\nu_{e}=\frac{1}{4}$ (c.f. the flat real and imaginary spectra in Fig. 2 and Supplementary Material [31]) is consistent with the fact that $C_{\mathrm{LC}}=0$ in these bands (not shown). In contrast, the $\Lambda=3,4$ states at $\nu_{e}=\frac{1}{4}$ exhibit a nonzero curvature in $\Gamma_{\bm{Q}}^{\Lambda}$ such that they sustain nonzero $\Omega$ in the LC.
Extracting band properties via optical reflectivity.— Finally, we discuss optical experiments to probe the aforementioned collective excitonic properties. More specifically, we illustrate that information about $\hat{H}^{\prime}$ can be obtained by measuring the reflection sector of the far-field scattering matrix for plane-wave incident light (near exciton resonances). For Bravais lattices (e.g., triangular and rectangular arrays), components of this matrix satisfy:
$S_{\tau,\tau^{\prime\prime}}(\bm{k})=\frac{3\pi c\gamma}{\omega_{\mathrm{ex}}}\sum_{\tau^{\prime}}g_{\tau,\tau^{\prime}}(k,\bm{k}_{||})D_{\tau^{\prime},\tau^{\prime\prime}}(ck,\bm{k}_{||}),$ (8)
where $\bm{k}_{||}$ is the in-plane component of $\bm{k}$ and $\tau$ ($\tau^{\prime}$) stands for the (circular) polarization of light here (which is locked to the valley index of electrons in TMDs [6]). $g_{\tau,\tau^{\prime}}(k,\bm{Q})=-\frac{i}{2\mathcal{A}_{\nu_{e}}}\frac{k^{2}\delta_{\tau,\tau^{\prime}}-(\bm{e}_{\tau}^{\ast}\cdot\bm{Q})(\bm{e}_{\tau^{\prime}}\cdot\bm{Q})}{k^{2}\sqrt{k^{2}-Q^{2}}}$ is the Fourier transform of the Green’s tensor [31]. $D_{\tau,\tau^{\prime}}(\omega,\bm{Q})=\langle\hat{x}_{\bm{Q},\tau}(\omega-\hat{H}^{\prime})^{-1}\hat{x}_{\bm{Q},\tau^{\prime}}^{\dagger}\rangle$ describes the collective response susceptibility of the exciton, evaluated in the vacuum state, with $\hat{x}_{\bm{Q},\tau}$ denoting the Fourier transform of $\hat{x}_{\bm{R},\tau}$. As shown in Eq. (8), measurements of $S_{\tau,\tau^{\prime}}(\bm{k})$ can help us determine $\hat{H}^{\prime}$ and the corresponding excitonic properties within the LC. In particular, in addition to standard spectral properties such as collective lineshifts and linewidths, we find that $C_{\mathrm{LC}}$ can be extracted from (half of) the winding number of phase of $S_{\tau,-\tau}$ with respect to the polar angle of $\bm{k}_{||}$, as shown in Fig. 4(b).
For non-Bravais lattices, however, $S_{\tau,\tau^{\prime}}(\bm{k})$ can only provide partial information about the exciton Hamiltonian because far-field measurements cannot resolve the sublattices. More specifically, $(\omega-\hat{H}^{\prime})^{-1}$ in $D_{\tau,\tau^{\prime}}(\omega,\bm{Q})$ has to be sublattice-symmetrized in this situation [31]. Therefore, measuring $S_{\tau,\tau^{\prime}}(\bm{k})$ generally cannot uniquely determine all the excitonic properties within the LC. One special case where such a scheme is still applicable is the honeycomb lattice, where sublattice symmetrization selects the $\Lambda=0,1$ doublet such that their $\Delta_{\bm{Q}}$, $\Gamma_{\bm{Q}}$, and $\Omega(\bm{Q})$ are still fully recoverable from reflection experiments  [^14] Note that radiation from $\Lambda=2,3$ states of the honeycomb lattice is suppressed and hence its contribution to the optical response is negligible for all practical purposes. A complete reconstruction of the excitonic band structures would require additional measurements beyond $S_{\tau,\tau^{\prime}}(\bm{k})$, which we leave for future work.
Outlook. — Our formalism can be further generalized to describe the collective behavior of moiré excitons involving other strong correlations to study their interplay. For instance, going beyond the single excitation subspace, nonlinearities inherent in these emitters can lead to a variety of interesting physical phenomena, including Dicke superradiance [53, 54], optical bistability [55], leaky condensation [56], and phase space filling [57, 58, 59]. Another intriguing problem is whether the spin correlations that emerge near half-filling $\nu_{e}=1$, such as the magnetic polaron effect [60, 61] and kinetic magnetism [62], could play a role in the cooperative excitonic properties.
In addition to WCs where the doped electrons only act as an effective lattice potential to the excitons, it is natural to ask about their collective behavior at general $\nu_{e}$. At fillings slightly away from the ones providing WCs, metastable frozen charge configurations could emerge and play the role of a random potential to emitters that breaks the translation symmetry [45]. We anticipate this randomness to suppress the degree of constructive interference, indicating weaker radiative decay.
Another possible outlook for moiré excitons is to simulate topological physics in two-dimensional dipolar spin systems [57]. In particular, their tunneling driven by dipole-dipole interaction could provide relatively flat collective bands with nontrivial Berry curvature within the light cone. These ingredients allow for the emergence of topological phases such as fractional Chern insulators and spin liquids [63, 64, 65]. We therefore anticipate these optical excitations in TMD bilayers to act as a platform for these phases of matter.
Acknowledgements. — We acknowledge A. Srivastava, S. Yelin and A. Asenjo-Garcia and D. Goncalves for useful discussions. Y.-Q. W. acknowledges the support from the JQI postdoctoral fellowship at the University of Maryland. D.E.C. acknowledges support from the European Union, under European Research Council grant agreement No 101002107 (NEWSPIN), FET-Open grant agreement No 899275 (DAALI) and EIC Pathfinder Grant No 101115420 (PANDA); the Government of Spain (Severo Ochoa Grant CEX2019-000910-S [MCIN/AEI/10.13039/501100011033]); QuantERA II project QuSiED, co-funded by the European Union Horizon 2020 research and innovation programme (No 101017733) and the Government of Spain (European Union NextGenerationEU/PRTR PCI2022-132945 funded by MCIN/AEI/10.13039/501100011033); Generalitat de Catalunya (CERCA program and AGAUR Project No. 2021 SGR 01442); Fundació Cellex, and Fundació Mir-Puig.