### v2

# Collective optical properties of moiré excitons

Abstract
We propose that excitons in moiré transition metal dichalcogenide bilayers offer a promising platform for investigating collective radiative properties. While some of these optical properties resemble those of cold atom arrays, moiré excitons extend to the deep subwavelength limit, beyond the reach of current optical lattice experiments. Remarkably, we show that the collective optical properties can be exploited to probe certain correlated electron states without requiring subwavelength spatial resolution. Specifically, we illustrate that the Wigner crystal states of electrons doped into these bilayers act as an emergent periodic potential for excitons. Moreover, the collective dissipative excitonic bands and their associated Berry curvature can reveal various charge orders that emerge at the corresponding electronic doping. Our study provides a promising pathway for future research on the interplay between collective effects and strong correlations involving moiré excitons.
Introduction. — The collective behavior of optical emitters in two-dimensional lattices driven by dipole-dipole interaction has generated considerable interest. Emitters in such arrays display a range of intriguing properties, including perfect reflection and transmission, dressed lineshift, enhanced and suppressed radiation, and even topological attributes such as collective Chern bands and protected edge states [1, 2, 3, 4]. Recently, this system has been experimentally realized using cold atoms in optical lattices [5], providing a promising avenue for the study of these coherent phenomena.
On the other hand, excitons (electron-hole bound states) in transition metal dichalcogenide (TMD) bilayers [6, 7, 8, 9, 10, 11] can form in a moiré lattice, providing an interesting platform to study collective optical properties. Moreover, the bilayers can host Wigner crystal (WC) states of doped electrons  [12, 13, 14, 15, 16, 17, 18, 19, 20], which can modify the excitonic properties.
In this work, we investigate the collective optical properties of moiré excitons in TMDs at incompressible states of doped electrons for various fractional fillings $\nu_{e}$  [^1] For clarity, the term “collective” in this work refers to the interference effects of states of an optical emitter from different sites [22, 23] rather than many-body physics of multiple emitters [24, 25, 26]. We focus on a weak excitation regime that yields dilute excitons, allowing us to neglect their electrostatic repulsion that manifests in high-power experiments [26, 20, 16, 18]. These excitations tunnel between empty moiré sites through dipole-dipole interaction, but cannot sit on top of a doped electron due to exciton-electron repulsion, as illustrated in Fig. 1(a). Therefore, WCs emerging at finite $\nu_{e}$ act as additional superpotentials, modifying the lattice potential experienced by a single exciton. Moreover, different $\nu_{e}$’s give rise to distinct WCs [see Fig. 1(b)], which induce distinctive excitonic dynamics with quantitatively different collective (spectral and topological) excitonic properties [27]. As such, those optical signatures serve as fingerprints of the corresponding WCs (see Table 1). Notably, the filling fractions $\nu_{e}$ can be easily accessed in gate-tunable experiments. Such tunability, combined with adjustable twist angles, further allows for the exploration of subwavelength emitter-array physics within different lattice configurations.
Figure 1: (a) Illustration of tunneling mediated by dipole-dipole interactions (green arrow) of an exciton (indicated by a pair of red and light blue dots) in a moiré potential (yellow) with period $a_{M}$. The potential can host a Wigner crystal state (WC) of doped electrons (dark blue dots), here a filling of $\nu_{e}=\frac{1}{3}$ is illustrated. Due to exciton-electron repulsion, the exciton cannot populate the sites occupied by electrons (indicated by the crossed-out dashed path). The remaining available sites form an emergent excitonic lattice. (b) WC at various $\nu_{e}$ in zero-twist WSe2/WS2 [14, 28]. The remaining empty sites provide a lattice for exciton tunneling shown in panel (a).
[…]
Emergent arrays $\nu_{e}$ Superradiant $\Lambda$ $C_{\mathrm{LC}}$
Triangular $0,\frac{2}{3},\frac{3}{4},\frac{6}{7}$ $0,1$ $-1,1$
Rectangular $\frac{1}{2}$ $0,1$ $0,0$
Honeycomb $\frac{1}{3}$ $0,1$ $-1,1$
Kagome $\frac{1}{4}$ $0,1,3,4$ $-1,1,1,-1$
Table 1: Emergent lattices and collective emission properties of interlayer excitons at various filling fractions ($\nu_{e}$), which form complementary lattices of charge orders in zero-twist WSe2/WS2 [14, 28]. The third column lists the collective band indices (in energy order) $\Lambda$ with enhanced radiative decay (within the light cone) compared to the bare dipole transition rate $\gamma$. $C_{\mathrm{LC}}$ denotes the corresponding Berry curvature, summed over the light cone, upon an out-of-plane magnetic field with Zeeman splitting $\mu_{B}B=20\gamma$.
Two-band model. — We consider the lowest conduction (CB) and the highest valence (VB) bands of two TMD monolayers. Each band is labeled by the valley pseudospin, which is locked to the real spin degrees of freedom for energy scales lower than the corresponding spin-orbit splittings [34]. Stacking these layers with a tunable twist angle or (for heterobilayers) lattice mismatch generally leads to interlayer coupling with an enlarged spatial periodicity $a_{M}$. We specifically consider heterobilayers [35], as the band offsets therein allow us to capture the effect of interlayer tunneling by emergent superlattice potentials, which split the original electronic dispersion into moiré bands  [^3] Such band misalignment also suppresses the topological effects favored by layer-degenerate states [78], which is not the main focus of our work.. In addition, we consider doping electrons into the first conduction moiré band and focus on the regime with dilute electron-hole pairs [6, 37], generated via weak optical excitation. Note that within such a dilute regime, the emission properties are captured by the single excitation subspace, in which the spectrum of a master equation is mathematically equivalent to the eigenspectrum of a non-Hermitian Hamiltonian [1, 2, 3]. While our formalism can be straightforwardly generalized, for simplicity, all (doped and optically excited) electrons are assumed to be in the same layer.
We start with the following two-band model in the absence of an external drive (see its microscopic origin in Ref. [38]):
^$\hat{\mathcal{H}}=\sum_{n=c,v}\sum_{\tau}\int d^{2}\bm{s}\hat{\psi}_{\tau}^{(n)\dagger}(\bm{s})h_{n}(\bm{s})\hat{\psi}_{\tau}^{(n)}(\bm{s})+\hat{\mathcal{V}}_{e}+\hat{\mathcal{V}}_{d}.$ (1)
[…]
$\hat{\mathcal{V}}_{d}=\frac{\omega_{p}^{2}}{c^{2}\epsilon_{0}}\int_{\bm{s},\bm{s}^{\prime}}\hat{\bm{P}}^{\dagger}(\bm{s})\cdot\bm{\mathcal{G}}\left(\omega_{p},\bm{s}-\bm{s}^{\prime}\right)\cdot\hat{\bm{P}}(\bm{s}^{\prime}),$ (2)
where $c$ is the speed of light and $\epsilon_{0}$ is the vacuum electric permittivity  [^6] Dielectric screening is neglected at the target frequency for simplicity.. $\hat{\bm{P}}^{\dagger}(\bm{s})=\sum_{\tau}\bm{d}_{\tau}^{cv}\hat{\psi}_{\tau}^{(c)\dagger}(\bm{s})\hat{\psi}_{\tau}^{(v)\dagger}(\bm{s})$ is the pair creation operator  [^7] R-stacked bilayers are considered such that the two charges have the same $\tau$ [79]., where $\bm{d}_{\tau}^{cv}$ is the transition dipole matrix element between the CB and VB at valley $\tau$. $\bm{\mathcal{G}}\left(\omega_{p},\bm{s}-\bm{s}^{\prime}\right)$ is the dyadic Green’s tensor evaluated at the frequency of the target pair state $\omega_{p}$ [38]. In the following, we use the two-band model $\hat{\mathcal{H}}$ to construct the superlattice Hamiltonian for optical excitations in the lowest energy manifold and doped electrons.
Zero doping. — At $\nu_{e}=0$, optical excitations from Eq. (1) are given by excitons. To capture the lowest composite particle, we employ the tight-binding approximation to the two-particle Hamiltonian operator from $\hat{\mathcal{H}}$  [^8] In the tight-binding regime, both the center-of-mass spatial fluctuations of the electron-hole pair from the moiré potential minima and the electron-hole separation are much smaller than the moiré period., whose eigenfunctions in the tight-binding limit can be approximated by moiré-Wannier orbitals [38]. The lowest energy orbital at each supersite $\bm{R}$, denoted as $w_{\bm{R}}(\bm{s}_{c},\bm{s}_{v})$ with $\bm{s}_{c}$ and $\bm{s}_{v}$ being the coordinates of the electron and hole, respectively, defines the corresponding exciton creation operator:
$\hat{x}_{\bm{R},\tau}^{\dagger}=\int d^{2}\bm{s}_{c}d^{2}\bm{s}_{v}w_{\bm{R}}(\bm{s}_{c},\bm{s}_{v})\hat{\psi}_{\tau}^{(c)\dagger}(\bm{s}_{c})\hat{\psi}_{\tau}^{(v)\dagger}(\bm{s}_{v}).$ (3)
Projecting $\hat{\mathcal{H}}$ onto these single-exciton basis states, the superlattice Hamiltonian becomes  [^9] Here, we drop the dipole-dipole interaction from the small out-of-plane transition dipole (not the permanent dipole) of interlayer excitons [51], which can be further suppressed via hybridization with intralayer excitons [80].:
$\hat{H}=\left(\omega_{\mathrm{ex}}-\frac{i\gamma}{2}\right)\hat{n}_{\mathrm{ex}}-\sum_{\bm{R}\neq\bm{R}^{\prime}}\sum_{\tau,\tau^{\prime}}t_{\bm{R},\bm{R}^{\prime}}^{\tau,\tau^{\prime}}\hat{x}_{\bm{R},\tau}^{\dagger}\hat{x}_{\bm{R}^{\prime},\tau^{\prime}},$ (4)
where $\hat{n}_{\mathrm{ex}}=\sum_{\bm{R},\tau}\hat{x}_{\bm{R},\tau}^{\dagger}\hat{x}_{\bm{R},\tau}$. $\omega_{\mathrm{ex}}$ and $\gamma$ denote the exciton frequency and decay rate (we set $\hbar=1$ hereafter), respectively, in a unit supercell problem, whereas the true excitation spectrum in the superlattice is renormalized by the tunneling:
$t_{\bm{R},\bm{R}^{\prime}}^{\tau,\tau^{\prime}}=t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}\delta_{\tau,\tau^{\prime}}-\frac{\omega_{\mathrm{ex}}^{2}|d|^{2}}{c^{2}\epsilon_{0}}\bm{e}_{\tau}^{\ast}\cdot\bm{\mathcal{G}}^{\mathrm{ex}}\left(\bm{R}-\bm{R}^{\prime}\right)\cdot\bm{e}_{\tau^{\prime}},$ (5)
which incorporates the sector from charge dynamics (denoted as $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}$) and the dipole-dipole interaction from the in-plane components of $\bm{d}_{\tau}^{cv}$  [^10] Similar interactions between dipoles within the same supercell has already been incorporated in $\omega_{\mathrm{ex}}$ and $\gamma$., denoted as $d\bm{e}_{\tau}^{\ast}$ with $\bm{e}_{\tau}=\frac{\bm{e}_{x}+i\tau\bm{e}_{y}}{\sqrt{2}}$ ($\bm{e}_{x}$ and $\bm{e}_{y}$ are in-plane unit vectors) [50, 51]. Here, $\bm{\mathcal{G}}^{\mathrm{ex}}\left(\bm{R}-\bm{R}^{\prime}\right)=\int_{\bm{s},\bm{s}^{\prime}}w_{\bm{R}}^{\ast}(\bm{s})\bm{\mathcal{G}}\left(\omega_{\mathrm{ex}};\bm{s}-\bm{s}^{\prime}\right)w_{\bm{R}^{\prime}}(\bm{s}^{\prime})$ where $w_{\bm{R}}(\bm{s})\equiv w_{\bm{R}}(\bm{s},\bm{s})$ is approximated as a Gaussian with width $a_{W}$ for simplicity. We refer to the supplementary material for details of $\hat{H}$ [38].
Finite doping with charge order. — We can generalize Eq. (4) to finite $\nu_{e}$, where doped electrons form WCs  [^11] The collective excitonic properties for hole-doped cases are qualitatively similar to electron-doped ones.. These charge orders could emerge from a generalized Hubbard model [53, 54], which naturally appears by projecting Eq. (1) onto the first moiré band of doped charges [38]. Importantly, since the exciton satisfies the tight-binding criterion, its hole remains bound to its original electron, preventing recombination with doped charges [55]. Further assuming that the electronic states being stable against the dynamics of dilute optical excitations, we can treat the WC as a spatially periodic detuning, consistent with experimental signatures [32, 33]. This detuning is characterized by the gap between $\omega_{\mathrm{ex}}$ and the energy of lowest three-body state (one doped electron together with the excited electron-hole pair) within a supercell, denoted as $\omega_{t}$. In particular, in the regime $|\omega_{t}-\omega_{\mathrm{ex}}|\gg|t_{\bm{R},\bm{R}^{\prime}}^{\tau,\tau^{\prime}}|$, the low-energy optical excitations are either within the $\sim\omega_{\mathrm{ex}}$ or $\sim\omega_{t}$ manifold, depending on the sign of $\omega_{t}-\omega_{\mathrm{ex}}$. We focus specifically on the case $\omega_{\mathrm{ex}}\ll\omega_{t}$, which is realizable by interlayer excitons [38, 20], to study the collective behavior of these emitters (three-body states with $\omega_{\mathrm{ex}}\gg\omega_{t}$ correspond to trions [56], which is not the focus of this work). In this subspace, these excitons are still described by the Hamiltonian Eq. (4121212Note that the energy integrals are different for the excitations in the $\sim\omega_{\mathrm{ex}}$ and $\sim\omega_{t}$ manifolds, due to their distinction in wavefunctions, except that now $\{\bm{R}\}$ lies in the emergent lattices, which are complementary lattices of the WCs. Accordingly, different emergent lattices can be realized simply by accessing distinct WCs in moiré TMD bilayers, which can be tuned by the electron filling fraction $\nu_{e}$ (doped electrons per supercell) [53, 14, 28]. We list a few possible emergent lattices observed in WSe2/WS2 in Table 1.
Figure 2: Collective excitonic lineshifts $\Delta_{\bm{Q}}$ (solid lines) and linewidths $\Gamma_{\bm{Q}}$ (dashed) emerging from charge-ordered zero-twist WSe2/WS2 with electron fillings (a) $\nu_{e}=0$ and (b) $\nu_{e}=\frac{1}{3}$. The vertical axes are displayed in units of $\gamma N_{\nu_{e}}$, with $N_{\nu_{e}}$ defined in Eq. (6). The horizontal axes represent the Bloch momentum $\bm{Q}$, which follows a piecewise-linear path through high-symmetry points in the Brillouin zone, as indicated by the hexagon in the inset of (a). Momenta within the light cone are indicated by the green shaded area (size enlarged for clarity). Different colors label distinct single-particle exciton bands. The parameters used are: $a_{M}=8.25$nm, $\omega_{\mathrm{ex}}=1.55$eV, and $a_{W}=2$nm [38, 26].
Collective bands. — We proceed to study the eigenspectrum of Eq. (4). To capture the relevant physics, we set $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}=0$ in Eq. (5) throughout this work  [^13] Neglecting $t_{\bm{R},\bm{R}^{\prime}}^{\mathrm{int}}$ is a valid approximation at large $a_{M}$ because it decays exponentially with $|\bm{R}-\bm{R}|/a_{M}$, whereas the dipole-dipole interaction only scales as a power law.. Diagonalization of the Hamiltonian Eq. (4) yields the emitter spectrum $\omega_{\mathrm{ex}}+\Delta_{\bm{Q}}-\frac{i}{2}\Gamma_{\bm{Q}}$ characterized by the center-of-mass Bloch momentum $\bm{Q}$. Note that, if the excitonic states at all $\bm{R}$ constructively interfere, $\Gamma_{\bm{Q}}$ should scale with $\gamma N_{\nu_{e}}(\theta)$ [1], where $N_{\nu_{e}}(\theta)$ is the ratio between the squared exciton wavelength and the unit cell area of the effective excitonic lattice at twisting angle $\theta$:
$N_{\nu_{e}}(\theta)=(1+\theta^{2}/\delta^{2})N_{\nu_{e}},\quad N_{\nu_{e}}=\lambda_{\mathrm{ex}}^{2}/\mathcal{A}_{\nu_{e}}.$ (6)
Here $\mathcal{A}_{\nu_{e}}$ denotes the emergent unit cell area at electron filling $\nu_{e}$ and zero twist and $\delta$ defines the lattice mismatch between the two monolayers [26]. We present the collective spectrum in units of $\gamma N_{\nu_{e}}(\theta)$ hereafter to indicate the extent of constructive interference of the exciton eigenstates.
In Fig. 2, we illustrate the collective lineshifts $\Delta_{\bm{Q}}$ and linewidths $\Gamma_{\bm{Q}}$ for the untwisted WSe2/WS2 bilayer at $\nu_{e}=0$ and $\frac{1}{3}$, corresponding to triangular and honeycomb lattices, respectively. The cases with Kagome and rectangular lattices can be found in the supplementary materials [38]. For all $\nu_{e}$ of interest, both quantities vary on scales of $N_{\nu_{e}}\gamma$ with $N_{\nu_{e}}\sim 10^{4}$. Qualitatively, this is because emissions from supersites much closer than the exciton wavelength $\lambda_{\mathrm{ex}}$ can constructively interfere.
The collective behavior is qualitatively distinct within and outside of the LC. For all $\nu_{e}$, $\Gamma_{\bm{Q}}$ is significantly greater than $\gamma$ within the LC but is suppressed outside, consistent with the fact that only emitters in the LC couple to light and radiate due to momentum conservation. Note also that such a large decay rate in the LC indicates that an emitter tends to radiate before it hops to other supersites, which is also reflected by the relatively flat $\Delta_{\bm{Q}}$ (with respect to $\bm{Q}$) therein.
Figure 3: Dependence of collective lineshifts and linewidths at $\bm{Q}=0$ (a) on $\nu_{e}$ for all states at zero twist and (b) on twisting angle $\theta$ and $\nu_{e}$, whose effects manifest in the combined factor $N_{\nu_{e}}(\theta)$, see Eq. (6). Dashed lines in (b) indicate linear fits with slope $1$. Here we set $\delta=0.04$ [26], and $a_{W}(\theta)=(1+\theta^{2}/\delta^{2})^{-\frac{1}{4}}a_{W}(0)$ [59]; all other parameters are the same as in Fig. 2.
Aside from these common properties, collective bands exhibit several distinct features at different $\nu_{e}$. We specifically illustrate how $\Delta_{0}$ and $\Gamma_{0}$ vary with $\nu_{e}$ in Fig. 3(a), as these quantities are directly accessible via reflection measurements at normal incidence [1]. First, different emergent lattices generally yield distinct numbers of bright levels, which serves as the most direct signature to identify certain WCs. For instance, two bright lines with equal and different $\Gamma_{0}$ can be associated to rectangular ($\nu_{e}=\frac{1}{2}$) and Kagome ($\nu_{e}=\frac{1}{4}$) lattices, respectively. In contrast, triangular and honeycomb lattices cannot be differentiated by bright-level counting.
Another optical signature of the WCs is the linewidth $\Gamma_{0}$ of the lowest state at each $\nu_{e}$, which is typically the fastest radiative rate. By comparing their values (which overall decreases with $\nu_{e}$), it is possible to determine whether the corresponding WCs are formed in experiments  [^14] While a similar comparison can be made for $\Delta_{0}$, we anticipate that additional factors beyond the scope of this study, including long-range exciton-electron interactions, may also influence the exciton energy. In spectral measurements, isolating $\Delta_{0}$ from those factors could be challenging..
Finally, these spectral properties can also be tuned via twisting angle $\theta$. Fig. 3(b) illustrates the dependence of $\Gamma_{0}$ on $N_{\nu_{e}}(\theta)$. Crucially, for the same type of lattice geometry, the collective linewidths scale inversely with the emergent unit cell area, which can be confirmed by the analytic expression of the Green’s tensor at $\bm{Q}=0$. These scalings provide an indirect probe for the emergent lattice geometry: for instance, all triangular arrays (e.g., $\nu_{e}=0,\frac{2}{3}$) belong to the same line in Fig. 3(b). Therefore, the intercept in this logarithmic plot serves as another fingerprint of the WCs.
Berry curvature. — Topological features can also emerge in these collective excitonic bands. To see this, we introduce a magnetic field to break time-reversal symmetry and compute the total Berry curvature within the LC, $C_{\mathrm{LC}}$, for (directly) optically accessible states, as summarized in Table 1. Excitonic bands arising from triangular, honeycomb, and Kagome effective lattices yield $C_{\mathrm{LC}}=\pm 1$, whereas those from rectangular lattices remain topologically trivial. This behavior originates from the phase winding of the light-mediated coupling between valley doublets, which effectively manifests only in lattices with $C_{3}$ rotational symmetry, rather than from topological electronic bands [61].
Finally, we propose a far-field reflection experiment to detect $C_{\mathrm{LC}}$, where the scattering matrix element extracted from cross-circularly polarized input and output fields, $S_{-+}$, is measured. Specifically, its momentum dependence, controlled by incidence direction, possesses difference number of nodes for excitonic bands with different topological features. We leave the details to the End Matter.
Outlook. — Our formalism can be further generalized to describe the collective behavior of moiré excitons involving other strong correlations to study their interplay [62]. For instance, going beyond the single excitation subspace [63], nonlinearities inherent in these emitters can lead to a variety of interesting physical phenomena, including Dicke superradiance [64, 65, 66], optical bistability [67], leaky condensation [68], and phase space filling [69, 70, 71]. Another intriguing problem is whether the spin correlations that emerge near half-filling $\nu_{e}=1$, such as the magnetic polaron effect [72, 73] and kinetic magnetism [56], could play a role in the cooperative excitonic properties.
In addition to WCs where the doped electrons only act as an effective lattice potential to the excitons, it is natural to ask about their collective behavior at general $\nu_{e}$. At fillings slightly away from the ones providing WCs, metastable frozen charge configurations could emerge and play the role of a random potential to emitters that breaks the translation symmetry [54]. We anticipate this randomness to suppress the degree of constructive interference, indicating weaker radiative decay.
Another possible outlook for moiré excitons is to simulate topological physics in two-dimensional dipolar spin systems [69]. In particular, their tunneling driven by dipole-dipole interaction could provide relatively flat collective bands with nontrivial Berry curvature within the light cone. These ingredients allow for the emergence of topological phases such as fractional Chern insulators and spin liquids [74, 75, 76]. We therefore anticipate these optical excitations in TMD bilayers to act as a platform for these phases of matter.
Acknowledgements. — We acknowledge A. Srivastava, S. Yelin and A. Asenjo-Garcia and D. Goncalves for useful discussions. The work at Maryland was supported by W911NF2010232, MURI FA9550-19-1-0399, DARPA HR00112530313 and Simons and Minta Martin Foundations. Y.-X.W. acknowledges support from a QuICS Hartree Postdoctoral Fellowship. Y.-Q. W. acknowledges the support from the JQI postdoctoral fellowship at the University of Maryland. D.E.C. acknowledges support from the European Union, under European Research Council grant agreement No 101002107 (NEWSPIN), FET-Open grant agreement No 899275 (DAALI) and EIC Pathfinder Grant No 101115420 (PANDA); the Government of Spain (Severo Ochoa Grant CEX2019-000910-S [MCIN/AEI/10.13039/501100011033]); QuantERA II project QuSiED, co-funded by the European Union Horizon 2020 research and innovation programme (No 101017733) and the Government of Spain (European Union NextGenerationEU/PRTR PCI2022-132945 funded by MCIN/AEI/10.13039/501100011033); Generalitat de Catalunya (CERCA program and AGAUR Project No. 2021 SGR 01442); Fundació Cellex, and Fundació Mir-Puig. All authors acknowledge the hospitality of the Kavli Institute for Theoretical Physics (KITP) supported by grant NSF PHY-1748958.
Appendix A End Matter
Appendix A: Details of Berry curvature. — Incorporating an out-of-plane external magnetic field $B$ adds a Zeeman splitting between the valley-degenerate doublets to Eq. (4). We accordingly proceed with the Hamiltonian $\hat{H}\to\hat{H}^{\prime}\equiv\hat{H}+\mu_{B}B\sum_{\tau=\pm,\bm{R}}\tau\hat{x}_{\bm{R},\tau}^{\dagger}\hat{x}_{\bm{R},\tau}$, where $\mu_{B}$ denotes the magnetic dipole. The Berry curvature $\Omega({\bm{Q}})=i\nabla_{\bm{Q}}\times\langle\bm{Q}|\nabla_{\bm{Q}}|\bm{Q}\rangle$ is computed for each band at several $\nu_{e}$, where $|\bm{Q}\rangle$ are right eigenstates of $\hat{H}^{\prime}$. Note that without loss of generality, the right eigenvectors of the non-Hermitian Hamiltonian are chosen to define the Berry curvature [81, 82] (a more rigorous study of non-Hermitian topology is left for future work). We specifically focus on $\Omega({\bm{Q}})$ of (directly) optically accessible states (i.e., within the LC) and compute the numerical results for $\nu_{e}=0$ and $\frac{1}{2}$ [see Fig. 4(a)], evaluated under a standard momentum discretization scheme [83].
At $\nu_{e}=0$, both collective excitonic states (labeled with respect to their energies by $\Lambda=0,1$) exhibit nontrivial and opposite $\Omega({\bm{Q}})$ centered at $\bm{Q}=0$. Notably, the summation of $\Omega({\bm{Q}})$ in LC, $C_{\mathrm{LC}}$, for these bands takes integer values $\mp 1$. As briefly mentioned in the main text, topological features here originates from the phase winding of $\bm{e}_{\tau}^{\ast}\cdot\bm{\mathcal{G}}^{\mathrm{ex}}\left(\bm{R}-\bm{R}^{\prime}\right)\cdot\bm{e}_{-\tau}$, which appears as the off-diagonal terms in the effective low momentum approximation of $\hat{H}^{\prime}$, denoted as $\hat{h}(\bm{Q})$ [38]. With $\bm{Q}=Q(\cos\phi\bm{e}_{x}+\sin\phi\bm{e}_{y})$, where $\phi$ is the polar angle of $\bm{Q}$, we have:
$\hat{h}(\bm{Q})\simeq\omega_{\mathrm{ex}}+\Delta_{0}-\frac{i\Gamma_{0}}{2}+\begin{bmatrix}\mu_{B}B&iJQ^{2}e^{-2i\phi}\\ iJQ^{2}e^{2i\phi}&-\mu_{B}B\end{bmatrix},$ (7)
where $J$ is a (generally complex) coefficient characterizing the band curvature around $\bm{Q}=0$. We find that Eq. (7) can approximately capture $\Omega(\bm{Q})$, particularly the sharp feature at $Q=\sqrt{|\mu_{B}B/J|}$.
In contrast, $C_{\mathrm{LC}}$ from generic arrays (see table 1) may be trivial if the low-$Q$ model of the target band is different from Eq. (7). For instance, both collective excitonic bands from the rectangular lattice at $\nu_{e}=\frac{1}{2}$ give zero Berry curvature, and hence $C_{\mathrm{LC}}=0$. This is due to the absence of $C_{3}$ rotational symmetry, leading to nonzero off-diagonal terms in the corresponding Hamiltonian at $Q=0$, which indicates a gap separating two topologically trivial bands [38].
Similar arguments also apply to Berry curvatures of higher collective states. More specifically, the suppressed $J$ for $\Lambda=2,3$ at $\nu_{e}=\frac{1}{3}$ and $\Lambda=2,5$ at $\nu_{e}=\frac{1}{4}$ (c.f. the flat real and imaginary spectra in Fig. 2 and Supplementary Material [38]) is consistent with the fact that $C_{\mathrm{LC}}=0$ in these bands (not shown). In contrast, the $\Lambda=3,4$ states at $\nu_{e}=\frac{1}{4}$ exhibit a nonzero curvature in $\Gamma_{\bm{Q}}^{\Lambda}$ such that they sustain nonzero $\Omega$ in the LC.
Figure 4: Comparison between properties of topological and non-topological collective bands from $\nu_{e}=0$ and $\nu_{e}=\frac{1}{2}$, respectively, with Zeeman splitting $\mu_{B}B=20\gamma$. (a) Dependence of Berry curvature ($\Omega$) on Bloch momentum magnitude $Q$ for the two bands at each $\nu_{e}$ (labeled by colors and $\Lambda$), with solid and dashed lines indicating results from the full model $\hat{H}^{\prime}$ and the low-momentum model Eq. (7), respectively. The vertical axis is in units of $J/\gamma$, where $J$ is fit from $\Gamma_{\bm{Q}}$ in Fig. 2(a) using a momentum sample within $Q\leq{\pi}/{25\lambda_{\mathrm{ex}}}$. The gray vertical line indicates the momentum satisfying $JQ^{2}=\mu_{B}B$. The inset shows that the dependence of $\Omega$ on the polar angle of $\bm{Q}$ is negligible. (b) The phase of $S_{-+}$ at $k=\frac{2\pi}{\lambda_{\mathrm{ex}}}$ and $\bm{k}_{||}=k_{x}\bm{e}_{x}+k_{y}\bm{e}_{y}$ for each $\nu_{e}$. All other parameters are the same as in Fig. 2.
Appendix B: Extracting band properties via optical reflectivity. — In this appendix, we discuss optical experiments to probe the aforementioned collective excitonic properties. The numerical results of Berry curvature for $\nu_{e}=0$ and $\frac{1}{2}$ are shown in Fig. 4(a), and their manifestation in optical measurements is illustrated in Fig. 4(b). Here, far-field reflection experiment with plane-wave incident light near exciton resonances is considered, and the target physical observables are the amplitude and phase of the scattered field, which can be obtained via standard homodyne technique [29]. Below we illustrate that information about $\hat{H}^{\prime}$ can be inferred from the scattering matrix extracted from such measurement. For Bravais lattices (e.g., triangular and rectangular arrays), components of this matrix satisfy:
$S_{\tau,\tau^{\prime\prime}}(\bm{k})=\frac{3\pi c\gamma}{\omega_{\mathrm{ex}}}\sum_{\tau^{\prime}}g_{\tau,\tau^{\prime}}(k,\bm{k}_{||})D_{\tau^{\prime},\tau^{\prime\prime}}(ck,\bm{k}_{||}),$ (8)
where $\bm{k}_{||}$ is the in-plane component of $\bm{k}$ and $\tau$ ($\tau^{\prime}$) stands for the (circular) polarization of light here (which is locked to the valley index of electrons in TMDs [6]). $g_{\tau,\tau^{\prime}}(k,\bm{Q})=-\frac{i}{2\mathcal{A}_{\nu_{e}}}\frac{k^{2}\delta_{\tau,\tau^{\prime}}-(\bm{e}_{\tau}^{\ast}\cdot\bm{Q})(\bm{e}_{\tau^{\prime}}\cdot\bm{Q})}{k^{2}\sqrt{k^{2}-Q^{2}}}$ is the Fourier transform of the Green’s tensor [38]. $D_{\tau,\tau^{\prime}}(\omega,\bm{Q})=\langle\hat{x}_{\bm{Q},\tau}(\omega-\hat{H}^{\prime})^{-1}\hat{x}_{\bm{Q},\tau^{\prime}}^{\dagger}\rangle$ describes the collective response susceptibility of the exciton, evaluated in the vacuum state, with $\hat{x}_{\bm{Q},\tau}$ denoting the Fourier transform of $\hat{x}_{\bm{R},\tau}$. Once $\hat{H}^{\prime}$ is reconstructed from $S_{\tau,\tau^{\prime}}(\bm{k})$ via Eq. (8), the corresponding spectral and topological excitonic properties within the LC could be determined. In particular, $C_{\mathrm{LC}}$ can be extracted from (half of) the winding number of phase of $S_{\tau,-\tau}$ with respect to the polar angle of $\bm{k}_{||}$, as shown in Fig. 4(b).
For non-Bravais lattices, however, $S_{\tau,\tau^{\prime}}(\bm{k})$ can only provide partial information about the exciton Hamiltonian because far-field measurements cannot resolve the sublattices. More specifically, $(\omega-\hat{H}^{\prime})^{-1}$ in $D_{\tau,\tau^{\prime}}(\omega,\bm{Q})$ has to be sublattice-symmetrized in this situation [38]. Therefore, measuring $S_{\tau,\tau^{\prime}}(\bm{k})$ generally cannot uniquely determine all the excitonic properties within the LC. One special case where such a scheme is still applicable is the honeycomb lattice, where sublattice symmetrization selects the $\Lambda=0,1$ doublet such that their $\Delta_{\bm{Q}}$, $\Gamma_{\bm{Q}}$, and $\Omega(\bm{Q})$ are still fully recoverable from reflection experiments  [^15] Note that radiation from $\Lambda=2,3$ states of the honeycomb lattice is suppressed and hence its contribution to the optical response is negligible for all practical purposes. A complete reconstruction of the excitonic band structures would require additional measurements beyond $S_{\tau,\tau^{\prime}}(\bm{k})$, which we leave for future work.